# Improving Ab-Initio Cryo-EM Reconstruction with Semi-Amortized Pose Inference

### v1

@@ Abstract
Cryo-Electron Microscopy (cryo-EM) is an increasingly popular experimental technique for estimating the 3D structure of macromolecular complexes such as proteins based on 2D images. These images are notoriously noisy, and the pose of the structure in each image is unknown a priori. Ab-initio 3D reconstruction from 2D images entails estimating the pose in addition to the structure. In this work, we propose a new approach to this problem. We first adopt a multi-head architecture as a pose encoder to infer multiple plausible poses per-image in an amortized fashion. This approach mitigates the high uncertainty in pose estimation by encouraging exploration of pose space early in reconstruction. Once uncertainty is reduced, we refine poses in an auto-decoding fashion. In particular, we initialize with the most likely pose and iteratively update it for individual images using stochastic gradient descent (SGD). Through evaluation on synthetic datasets, we demonstrate that our method is able to handle multi-modal pose distributions during the amortized inference stage, while the later, more flexible stage of direct pose optimization yields faster and more accurate convergence of poses compared to baselines. Finally, on experimental data, we show that our approach is faster than state-of-the-art cryoAI and achieves higher-resolution reconstruction.
@@ 1 Introduction
Figure 1: Our method for ab-initio reconstruction consists of two stages: (i) an auto-encoding stage where an image encoder equipped with multiple heads maps the input image to the pose candidate set $\{R_{1},\dots R_{M}\}$, followed by computing projections by slicing through the volume decoder in Fourier space based on the pose set. These projections are compared with the input image and the one with the minimum error is used. (ii) An auto-decoding stage where pose parameters in axis-angle representation are stored for all images. The same volume decoder is used to obtain projections, and the reconstruction loss is computed for a single projection.
@@
Single particle cryo-EM has gained popularity among structural biologists as a powerful experimental method for determining the 3D structure of macromolecular complexes, such as proteins and viruses, a key to unlocking our understanding of biological function at the scale of the cell. Thanks to pioneering advances in hardware and data processing techniques, cryo-EM has enabled reconstruction of challenging structures at atomic or near-atomic resolution [1]. During a cryo-EM experiment, a particle stack of $10^{4-7}$ images of a target bio-molecule are acquired using an electron microscope, from which the goal is to reconstruct the unknown 3D density map [2].
@@
This ab-initio reconstruction task presents some challenges. First, the pose (i.e., orientation and position) of the particle in each observation is unknown, thus the pose must be estimated for each image. Second, the images have a low signal-to-noise ratio (SNR) to avoid electron damage during imaging (e.g., see as shown in Fig. 1). The low SNR obscures high-resolution image detail, hindering pose and structure estimation. Third, bio-molecules are typically non-rigid and exhibit structural variations within a sample. Hence, for such heterogeneous datasets, it is crucial to account for the conformational variability in order to obtain high-resolution reconstruction [3, 4, 5, 6].
@@
Recent state of the art methods development in cryo-EM have focused on deep learning. CryoDRGN [5] proposed an image encoder-volume decoder architecture to model continuous heterogeneity, but the poses were assumed to be given. CryoDRGNv2 [7] introduced hierarchical pose search, comprising grid search followed by a form of branch-and-bound (BnB) optimization, akin to cryoSPARC [8]. More recently, cryoPoseNet [9] and cryoAI [10] introduce amortization to pose inference, to avoid the expense of orientation matching for each image. They employ a convolutional neural network (CNN) to map an input image to a pose estimate. While more efficient, amortized inference could fail to represent multi-modal posterior pose distributions that usually occur early in reconstruction. As such, compared to previous methods like cryoSPARC [8] and cryoDRGNv2 [7] that consider many pose candidates, amortized methods sometimes fail to identify the correct mode. Moreover, amortization confines the pose search to a global parametric function (the encoder) which can lead to slow convergence of poses.
@@
Here, we introduce a new approach to ab-initio homogeneous reconstruction that is able to handle multi-modal pose distributions with a tailored encoder, and accelerates pose optimization with semi-amortization [11]. First, we perform amortized pose inference through a shared multi-choice encoder that maps each input image to multiple pose estimates (Fig. 1, top). In contrast to cryoPoseNet [9] and cryoAI [10], which use pose encoders that produce one or two estimates, we attach multiple pose predictor heads to a shared CNN feature extractor to predict multiple plausible poses for each image. By design, our multi-output encoder is able to account for pose uncertainty and encourage exploration of pose space during initial stages of reconstruction. For the 3D decoder, unlike cryoDRGN and cryoAI which use MLPs with computationally expensive feed-forward implicit networks, we adopt an explicit parameterization to further accelerate the reconstruction. To train the encoder-decoder architecture, we apply a “winner-takes-all” loss in which the 3D decoder is queried to obtain a 3D-to-2D projection for each predicted pose. Inspired by the loss function in Multi-Choice learning [12, 13, 14], we select the projection with the lowest image reconstruction error to determine the loss.
@@
During the course of optimization, as higher resolution details are resolved, the pose posterior becomes uni-modal; thus, the pose search can be narrowed down to a neighborhood containing the most likely mode. In this stage, we propose to transition from auto-encoding to auto-decoding (Fig. 1, bottom). In particular, for each image, we choose the pose with the lowest reconstruction loss, iteratively refine it with stochastic gradient descent (SGD), and continue pose optimization and reconstruction in an alternating fashion. This direct, per-image optimization procedure achieves faster convergence to more accurate poses than amortized inference alone, which relies on potentially sub-optimal predictions of the encoder as a parametric function.
@@
We evaluate our approach and compare it with two state-of-the-art methods, namely, cryoSPARC [8] and cryoAI [10], on synthetic and real experimental datasets. Using synthetic data, we validate that semi-amortized inference noticeably accelerates the convergence of poses and yields reconstructions at similar or higher resolutions. We also show that our multi-choice encoder empirically accounts for multiple modes in the pose posterior. Finally, we apply our method to an experimental dataset and obtain reconstruction outperforming cryoAI and competitive with cryoSPARC.
@@
  • We train the encoder coupled with an explicit volume decoder using a “winner-takes-all” loss that penalizes the pose hypothesis with least reconstruction error.
  • We introduce semi-amortization to pose inference which begins with amortized pose estimation by an encoder, followed by direct pose optimization for individual images.
  • In ab-initio reconstruction, our semi-amortized method is faster than amortized method of cryoAI and achieves higher-resolution on synthetic and experimental datasets.
@@ 2 Related work
Conventional homogeneous reconstruction techniques rely on common-lines [16, 17, 18] or projection mapping [19, 20] to select optimal poses. Other works  [21, 22] frame the reconstruction problem in the context of maximum a posteriori (MAP) estimation, and jointly reconstructs poses and structure via expectation maximization (EM). We compare our approach to cryoSPARC [8], a state-of-the-art method that uses stochastic gradient descent and a branch-and-bound search [23] for ab initio reconstruction and pose estimation. Like these methods, our auto-decoding stage directly optimizes pose of every image.
More recently, amortized inference techniques have been proposed for pose estimation [24, 9, 10, 15]. These techniques avoid explicit per-image pose optimization; instead, they train an auto-encoder or variational one [25] to associate each particle image with a predicted pose [9]. One challenge is that the auto-encoders can become stuck in local optima during training [9]. To address this issue, cryoAI [10] produces two pose estimates per image coupled with a symmetrized loss function that penalizes the best one. We build on this concept by adopting a multi-head neural architecture as the encoder to output multiple plausible pose candidates and avoid local optima.
@@
Inspired by scenarios where a set of hypotheses needs to be generated to account for uncertainty in the prediction task, MCL [12] was introduced in a supervised setup to learn multiple structured-outputs with SSVMs [26]. Their motivating question was: can we learn to produce a set of plausible hypotheses? To address this, they define an“oracle” loss in which only the most accurate output pays the penalty. This loss is minimized even if there is only a single accurate prediction in the set. The early follow-up work [27, 13] uses the same loss to learn a deep CNN ensemble composed of $M$ heads with a shared backbone network. Importantly, they show that the ensemble-mean loss hurts prediction diversity across different heads, while training with the “oracle“ loss yields specialized heads. Variations have since been proposed to mitigate hypothesis collapse or overconfidence issues in MCL by modifying the loss or applying learnable probabilistic scoring schemes [28, 29, 30, 14]. MCL has been used to mitigate the ambiguity in several tasks including image segmentation [31], optical-flow estimation [32], trajectory forecasting [33], human pose and shape estimation [14, 34]. In our work, we use the “oracle” loss in context of auto-encoder which supervises the pose encoder indirectly through projections provided by the decoder.
@@ 3 Problem Definition
The image formation model for cryo-EM is often well-approximated using the weak-phase object model [35]. This model assumes the 3D structure is an unknown density map, $V:\mathbb{R}^{3}\to\mathbb{R}_{\geq 0}$, represented under a canonical orientation. Cryo-EM images, $\{I_{i}\}_{i=1}^{N}$, are approximated as orthographic projections of the 3D map that are oriented and shifted by unknown rotation $R_{i}\in SO(3)$ and in-plane translation $t_{i}=(t_{x},t_{y})\in\mathbb{R}^{2}$. Formally,
@@
By the Fourier slice theorem [36], the Fourier transform of a projection is equal to a central slice through the density map’s 3D Fourier spectrum. Consequently,
@@
Under this model, given the structure $\hat{V}$, the negative log-likelihood of observing image $\hat{I}_{i}$ with noise variance $\sigma^{2}$ and poses $(R_{i},t_{i})$ is
@@
Ab-initio reconstruction methods [8, 9, 10, 7] solve jointly for the unknown structure $\hat{V}$ and poses $(R_{i},t_{i})$. They often follow an Expectation-Maximization (EM) [37, 21] procedure in which the E-step aligns images with the structure yielding pose estimates $(R_{i},t_{i})$, and then in the M-step the volume $\hat{V}$ is updated by minimizing the negative log-likelihood in Eq. 3. Since errors in pose estimates lead to blurry reconstructions, accurate pose estimates are crucial to finding high-resolution structures. As discussed above, poses are either optimized through search and projection matching [8, 7] or estimated by an encoder network [10, 9, 15].
@@ 4 Methodology
We propose a two-staged approach to ab-initio cryo-EM reconstruction by mixing auto-encoding and auto-decoding. In particular, we start with amortized inference (Fig. 1, auto-encoding) where a shared encoder, equipped with multiple heads, outputs a set of pose guesses. Once the pose posterior becomes less uncertain, we switch to direct pose optimization (Fig. 1, auto-decoding). The former enables handling the pose uncertainty early in reconstruction while the latter circumvents sub-optimality of encoder network leading to arguably more accurate poses and faster convergence. We couple our pose estimation module with an explicit volumetric decoder representing the 3D structure in Hartley space [5]. Our explicit model enables faster evaluation of projections compared to multiple passes through implicit neural representations [38, 39, 40]. The following sections discuss these components in detail.
@@ 4.1 Multi-choice encoder
With randomly initialized or a low-resolution reconstruction, the pose posterior contains multiple modes. Also, due to high levels of noise in images and near-symmetries in biological structures, pose estimation inherently involves high uncertainty. As a result, there exist several equally-plausible poses for each image, and naive optimization or search is prone to local minima.
To account for uncertainty in pose estimation, we build upon cryoAI [10] and extend the encoder to return multiple rotations as candidate poses. Formally, given the $i$-th image, $I_{i}\in\mathbb{R}^{H\times W}$, multiple rotations are computed as,
$F_{i}$ $=\textsc{VGG}(I_{i})$ (4)
$R_{i,j}$ $=H_{\theta_{j}}(F_{i}),\quad 1\leq j\leq M$ (5)
where a shared convolutional backbone based on the VGG16 architecture [41] extracts image-specific intermediate features, $F_{i}\in\mathbb{R}^{C\times H\times W}$. These features are then supplied to $M$ separate pose predictor heads, $H_{\theta_{j}},1\leq j\leq M$, yielding the pose set $\{R_{i,1},\dots,R_{i,M}\}$. Heads are composed of fully-connected layers with distinct learnable weights, while the backbone is shared across all heads. Using multiple heads facilitates pose exploration and reduces uncertainty in the pose estimation process. Moreover, as $\theta_{j}$ are randomly initialized, different heads are free to specialize on subsets of pose space so that they collectively produce a set of likely poses $\{R_{i,1},\dots,R_{i,M}\}$. In practice, we find that different heads specialize to non-overlapping regions of $SO(3)$ space (see Supp. B) and are thus complementary for accurate pose estimation across the entire dataset.
Inspired by multi-choice learning [12, 13, 14], we optimize the encoder-decoder using a “winner-takes-all” loss. For each image $I_{i}$ and predicted pose $R_{i,j}$, the negative log-likelihood, $\mathcal{L}_{i,j}$ in (Eq. 3), is computed, and the minimum is selected as the final loss for the corresponding image, i.e.,
@@
Minimizing this loss requires only one of the predicted poses to be accurate. CryoAI [10] can be viewed a special case of this formulation; it assigns two poses to each image as a consequence of input augmentation, and selects the best one with a symmetrized loss. In contrast, our approach augments the output of the encoder network with multiple heads, each providing a pose estimate.

### v2

# CryoSPIN: Improving Ab-Initio Cryo-EM Reconstruction with Semi-Amortized Pose Inference

@@ Abstract
Cryo-EM is an increasingly popular method for determining the atomic resolution 3D structure of macromolecular complexes (eg, proteins) from noisy 2D images captured by an electron microscope. The computational task is to reconstruct the 3D density of the particle, along with 3D pose of the particle in each 2D image, for which the posterior pose distribution is highly multi-modal. Recent developments in cryo-EM have focused on deep learning for which amortized inference has been used to predict pose. Here, we address key problems with this approach, and propose a new semi-amortized method, cryoSPIN, in which reconstruction begins with amortized inference and then switches to a form of auto-decoding to refine poses locally using stochastic gradient descent. Through evaluation on synthetic datasets, we demonstrate that cryoSPIN is able to handle multi-modal pose distributions during the amortized inference stage, while the later, more flexible stage of direct pose optimization yields faster and more accurate convergence of poses compared to baselines. On experimental data, we show that cryoSPIN outperforms the state-of-the-art cryoAI in speed and reconstruction quality. Project Webpage.
@@ 1 Introduction
Figure 1: CryoSPIN consists of two stages: (i) an auto-encoding stage where an image encoder equipped with multiple heads maps the input image to the pose candidate set $\{\phi_{1},\dots\phi_{M}\}$, followed by computing projections by slicing through the volume decoder in Fourier space based on the pose set. These projections are compared with the input image and the one with the minimum error is used. (ii) An auto-decoding stage where pose parameters in axis-angle representation are stored for all images. The same volume decoder is used to obtain projections, and the reconstruction loss is computed for a single projection.
@@
Single particle electron cryo-microscopy (cryo-EM) has gained popularity among structural biologists as a powerful experimental method for determining the 3D structure of macromolecular complexes, such as proteins and viruses, unlocking our understanding of biological function at the scale of the cell. Thanks to pioneering advances in hardware and data processing techniques, cryo-EM has enabled reconstruction of challenging structures at atomic or near-atomic resolution [1].
@@
During a cryo-EM experiment, $10^{4}$$10^{7}$ particle images, each containing an instance of the target bio-molecule, are acquired using a transmission electron microscope. From that particle stack, the goal is to reconstruct the unknown 3D density map [2]. This ab-initio reconstruction task presents some challenges. First, the 3D pose (orientation and position) of the particle in each image is unknown, and must be inferred along with the 3D structure. Second, low electron exposures are used to limit radiation damage to the particles (e.g., see Fig. 1). But this reduces the signal-to-noise ratio (SNR), obscuring high-resolution image detail and hindering pose and structure estimation. Third, bio-molecules are typically non-rigid and exhibit structural variations within a sample. Hence, for such heterogeneous datasets, it is crucial to account for the conformational variability in order to obtain high-resolution reconstruction [3, 4, 5, 6].
@@
The development of methods in cryo-EM has recently focused on deep learning. CryoDRGN [5] proposed an image encoder-volume decoder architecture to model continuous heterogeneity, but the poses were assumed to be known. CryoDRGNv2 [7] introduced hierarchical pose search, comprising grid search followed by a form of branch-and-bound (BnB) optimization, akin to cryoSPARC [8]. More recently, cryoPoseNet [9] and cryoAI [10] introduced amortized pose inference, to avoid the expense of orientation matching for each particle image. They employ a convolutional neural network (CNN) to map an input image to a pose estimate. Nevertheless, although efficient, amortized inference may not accurately represent multi-modal posterior pose distributions, which occur frequently early in reconstruction before the structure is estimated well. As such, compared to previous methods like cryoSPARC [8] and cryoDRGNv2 [7] that consider many pose candidates, amortized methods sometimes fail to identify the correct mode. Moreover, amortization confines the pose search to a global parametric function (the encoder) which can lead to slow convergence.
@@
Here, we introduce a new approach, called cryoSPIN, to ab-initio homogeneous reconstruction that is able to handle multi-modal pose distributions with a tailored encoder, and accelerates pose optimization with semi-amortization [11]. First, we perform amortized pose inference through a shared multi-choice encoder that maps each input image to multiple pose estimates (Fig. 1, top). In contrast to cryoPoseNet [9] and cryoAI [10], which use pose encoders that produce one or two estimates, we attach multiple pose predictor heads to a shared CNN feature extractor to predict multiple plausible poses for each image. By design, our multi-output encoder is able to account for pose uncertainty and encourage exploration of pose space during the initial stages of reconstruction. For the 3D decoder, unlike cryoDRGN and cryoAI that use MLPs with computationally expensive feed-forward implicit networks, we adopt an explicit parameterization to further accelerate the reconstruction. To train the encoder-decoder architecture, we apply a “winner-takes-all” loss in which the 3D decoder is queried to obtain a 3D-to-2D projection for each predicted pose. Inspired by the loss function in Multi-Choice Learning [12, 13, 14], we select the projection with the lowest image reconstruction error to determine the loss.
@@
During the course of optimization, as higher resolution details are resolved, the pose posterior becomes predominantly uni-modal. Thus, the pose search can be narrowed down to a neighborhood containing the most likely mode. In this stage, we propose to transition from auto-encoding to auto-decoding (Fig. 1, bottom). In particular, for each image, we choose the pose with the lowest reconstruction loss, iteratively refine it with stochastic gradient descent (SGD), and continue pose optimization and reconstruction in an alternating fashion. This direct, per-image optimization procedure achieves faster convergence to more accurate poses than amortized inference alone, which relies on potentially sub-optimal predictions of the encoder as a parametric function.
@@
We evaluate cryoSPIN in comparison to state-of-the-art methods, cryoSPARC [8], cryoAI [10], and cryoDRGN [7] on both synthetic and real experimental datasets. Using synthetic data, we find that semi-amortized inference noticeably accelerates the convergence of poses, and yields 3D reconstructions at similar or higher resolutions than baseline methods. We also show that our multi-choice encoder empirically accounts for multiple modes in the pose posterior. Finally, we apply cryoSPIN to an experimental dataset and obtain reconstructions with higher resolution than those from cryoAI, while matching the resolution of cryoSPARC reconstructions.
@@
  • We train the encoder coupled with an explicit volume decoder using a winner-take-all loss, which penalizes the pose hypothesis with least reconstruction error.
  • Beyond the architecture and objective, we introduce semi-amortized pose inference, which begins with feed-forward amortized pose inference, followed by direct, per-particle pose optimization.
  • In ab-initio reconstruction, our semi-amortized method, cryoSPIN, is faster than amortized method of cryoAI and achieves higher-resolution on synthetic and experimental datasets.
@@ 2 Related work
Early reconstruction techniques rely on common-lines [16, 17, 18] or projection mapping [19, 20] to select optimal poses. Other works  [21, 22] frame the reconstruction problem in the context of maximum a posteriori (MAP) estimation, and jointly reconstruct pose and structure via expectation maximization (EM). We compare our approach to cryoSPARC [8], a state-of-the-art method that uses stochastic gradient descent and a branch-and-bound search [23] for ab initio reconstruction and pose estimation. Like these methods, our auto-decoding stage directly optimizes pose of every image.
More recently, amortized inference techniques have been proposed for pose estimation [24, 9, 10, 15]. These techniques avoid explicit per-particle pose optimization; instead, they train an auto-encoder or variational one [25] to associate each particle image with a predicted pose [9]. One challenge is that the auto-encoders can become stuck in local optima during training [9]. To address this issue, cryoAI [10] produces two pose estimates per image coupled with a symmetrized loss function that penalizes the best one. We build on this concept by adopting a multi-head neural architecture as the encoder to output multiple plausible pose candidates and avoid local optima.
@@
Inspired by scenarios where a set of hypotheses needs to be generated to account for uncertainty in the prediction task, MCL [12] was introduced in a supervised setup to learn multiple structured-outputs with SSVMs [26]. Their motivating question was: can we learn to produce a set of plausible hypotheses? To address this, they define an“oracle” loss in which only the most accurate output pays the penalty. This loss is minimized even if there is only a single accurate prediction in the set. The early follow-up work [27, 13] uses the same loss to learn a deep CNN ensemble composed of $M$ heads with a shared backbone network. Importantly, they show that the ensemble-mean loss hurts prediction diversity across different heads, while training with the “oracle” loss yields specialized heads. Variations have since been proposed to mitigate hypothesis collapse or overconfidence issues in MCL by modifying the loss or applying learnable probabilistic scoring schemes [28, 29, 30, 14]. MCL has been used to mitigate the ambiguity in several tasks including image segmentation [31], optical-flow estimation [32], trajectory forecasting [33], human pose and shape estimation [14, 34]. In our work, we use the “oracle” loss in context of auto-encoder which supervises the pose encoder indirectly through projections provided by the decoder.
@@ 3 Problem Definition
Image formation in cryo-EM is well-approximated using the weak-phase object model [35]. Under this model, the structure of interest is an unknown 3D density map, $V:\mathbb{R}^{3}\to\mathbb{R}_{\geq 0}$, represented in some canonical coordinate frame. Cryo-EM images, $\{I_{i}\}_{i=1}^{N}$, are approximated as orthographic projections of the 3D map under some unknown coordinate transformation. We denote the unknown orientation by a rotation $R_{i}\in SO(3)$ and the unknown position in terms of an in-plane translation $t_{i}=(t_{x},t_{y})\in\mathbb{R}^{2}$. Formally,
@@
By the Fourier slice theorem [36], the Fourier transform of a projection is equal to a central slice through the density map’s 3D Fourier spectrum. That is,
@@
Under this model, given the structure $\hat{V}$, the negative log-likelihood of observing image $\hat{I}_{i}$ with noise variance $\sigma^{2}$ and pose $(R_{i},t_{i})$ is
@@
Ab-initio reconstruction methods [8, 9, 10, 7] solve jointly for the unknown structure $\hat{V}$ and per-particle poses $(R_{i},t_{i})$. They often follow an Expectation-Maximization (EM) [37, 21] procedure in which the E-step aligns images with the structure yielding pose estimates $(R_{i},t_{i})$, and then in the M-step the volume $\hat{V}$ is updated by minimizing the negative log-likelihood in Eq. 3. Since errors in pose estimates lead to blurry reconstructions, accurate pose estimates are crucial to finding high-resolution structures. As discussed above, poses are either optimized through search and projection matching [8, 7] or estimated by an encoder network [10, 9, 15].
@@ 4 Methodology
We introduce cryoSPIN, a two-staged approach to ab-initio cryo-EM reconstruction, combining auto-encoding and auto-decoding. We start with amortized inference (Fig. 1, auto-encoding) where a shared encoder equipped with multiple heads provide a set of pose guesses. Multiple guesses enable multimodal inference in the initial stages of reconstruction when the 3D structure is poorly resolved. As reconstruction improves, the pose posterior becomes less uncertain, at which point we switch to direct pose optimization (Fig. 1, auto-decoding). The former enables handling the pose uncertainty early in reconstruction while the latter circumvents sub-optimality of encoder network leading to arguably more accurate poses and faster convergence. We couple our pose estimation module with an explicit volumetric decoder representing the 3D structure in Hartley space [5]. Our explicit model enables faster evaluation of projections compared to multiple passes through implicit neural representations [38, 39, 40]. The following sections discuss these components in detail.
@@ 4.1 Multi-choice encoder
With randomly initialized or a low-resolution reconstruction, the pose posterior contains multiple modes. Also, due to high levels of noise in images and near-symmetries in biological structures, pose estimation inherently involves high uncertainty. As a result, there exist several plausible poses for each image, and naive optimization or search is prone to local minima.
To account for uncertainty in pose estimation, we build upon cryoAI [10] and extend the encoder to return multiple candidate poses $\phi$. Formally, given the image $I_{i}\in\mathbb{R}^{H\times W}$, $M$ pairs of rotations and translations are computed as,
$\phi_{i,j}\,=\,(R_{i,j},t_{i,j})\,=\,H_{\theta_{j}}(F_{i}),\quad 1\leq j\leq M$ (4)
where $R_{i,j}\in SO(3)$ and $t_{i,j}\in\mathbb{R}^{2}$ The image-specific intermediate features, $F_{i}=\textsc{VGG}(I_{i})\in\mathbb{R}^{C\times H\times W}$ are extracted by a shared convolutional backbone based on the VGG16 architecture [41], which are then supplied to $M$ separate pose predictor heads, $H_{\theta_{j}},1\leq j\leq M$, yielding poses $\phi_{i,j}$. The heads are composed of fully-connected layers with distinct learnable weights, while the backbone is shared across all heads. Using multiple heads facilitates pose exploration and reduces uncertainty in the pose estimation process. Moreover, as $\theta_{j}$ are randomly initialized, different heads are free to specialize on subsets of pose space so that they collectively produce a set of likely poses $\{\phi_{i,1},\dots,\phi_{i,M}\}$.
Inspired by multi-choice learning [12, 13, 14], we optimize the encoder-decoder using a “winner-take-all” loss. For each image $I_{i}$ and predicted pose $\phi_{i,j}$, the negative log-likelihood, $\mathcal{L}_{i,j}$ in (Eq. 3), is computed, and the minimum is selected as the final loss for the corresponding image, i.e.,
@@
Minimizing this loss requires only one of the predicted poses to be accurate. Interestingly, we find that this approach produces heads that specialize to somewhat disjoint regions of pose space (see Supp. C), consistent with previous work showing that it improves diversity in prediction tasks [13, 27]. This ’divide-and-conquer’ approach facilitates pose space exploration, which is crucial in order to avoid local minima during the early stages of reconstruction where uncertainty in the 3D structure is significant. CryoAI [10] can be viewed a special case of this formulation; it assigns two poses to each image as a consequence of input augmentation, and selects the best one with a symmetrized loss. In contrast, our approach augments the output of the encoder network with multiple heads, each providing a pose estimate.